The average height of 18-year-old males is 70 inches. A random sample of sixteen 18-year old males is selected. The average heights and standard deviation of heights came out to be 68.8 and 8 inches, respectively. At level of significance can we conclude that the average height has changed? a = .01

a) Set up test hypothesis

b) Write test statistics and sketch.

c) Apply classical or P-value method

d) Draw conclusion and interpret.
Answer #1

Solution:

Given:

Sample Size = n = 16

Sample mean = \bar{x} = 68.8 in.

Sample standard deviation = s = 8 in.

Population Mean = \mu = 70 in.

level of significance = 0.01

We have to test if  the average height has changed.

Part a) Set up test hypothesis

H_{0}: \mu = 70

H_{1}: \mu \neq 70

Part b) Write test statistics and sketch.

t=\frac{ \bar{x}-\mu }{s /\sqrt{n}}

t=\frac{ 68.8 - 70 }{ 8 /\sqrt{16 }}

t=\frac{ -1.2 }{ 8 / 4}

t=\frac{ -1.2 }{ 2 }

t = -0.600

t-0.600

Part c) Apply classical or P-value method

Using classical method: find t critical values:

df = n - 1 = 16 -1 = 15

Two tail area = level of significance = 0.01

t Table cum. prob one-tail two-tails t 50 t 995 75 0.25 ss 0.01 975 85 90 95 0.50 0.20 0.15 0.10 0.05 0.025 0.005 0.50 0.30 1

t critical values = \pm 2.947

Part d) Draw conclusion and interpret.

Decision Rule:

Reject null hypothesis H0, if absolute t test statistic value > t critical value = 2.947 , otherwise we fail to reject H0

since absolute t test statistic value = 0.600 < t critical value = 2.947 , we fail to reject H0.

Thus at 0.01 level of significance we do not have sufficient evidence to conclude that the average height has changed.